# Evaluating Human Alignment and Model Faithfulness of LLM Rationale

### v1

@@ Abstract
We study how well large language models (LLMs) explain their generations with rationales – a set of tokens extracted from the input texts that reflect the decision process of LLMs. We examine LLM rationales extracted with two methods: 1) attribution-based methods that use attention or gradients to locate important tokens, and 2) prompting-based methods that guide LLMs to extract rationales using prompts. Through extensive experiments, we show that prompting-based rationales align better with human-annotated rationales than attribution-based rationales, and demonstrate reasonable alignment with humans even when model performance is poor. We additionally find that the faithfulness limitations of prompting-based methods, which are identified in previous work, may be linked to their collapsed predictions. By fine-tuning these models on the corresponding datasets, both prompting and attribution methods demonstrate improved faithfulness. Our study sheds light on more rigorous and fair evaluations of LLM rationales, especially for prompting-based ones. [^1] Code and data will be released upon paper acceptance.
@@ 1 Introduction
The rise of large language models (LLMs) has significantly transformed the field of natural language processing (NLP) (Touvron et al., 2023; Team et al., 2023; OpenAI et al., 2024), enabling a wide range of applications from web question answering to complex reasoning tasks. However, they are not always reliable and usually cannot clearly explain their outputs (Ji et al., 2023), which limits the deployment of these models in high-stakes scenarios.
Rationales [^2] Also called self-explanation or extractive rationales in previous work (Huang et al., 2023a; Madsen et al., 2024)., i.e., tokens of the input text that are most influential to the models’ predictions, are widely studied in the NLP community prior to the era of LLMs to interpret model predictions (Lei et al., 2016; DeYoung et al., 2019; Wiegreffe and Pinter, 2019; Jacovi and Goldberg, 2020). For smaller and open-source models like BERT (Devlin et al., 2019), rationales are extracted with attribution-based methods like attention weights (Wiegreffe and Pinter, 2019) or gradients (Li et al., 2016). For decoder-only LLMs, besides attribution-based methods, rationales can also be extracted by leveraging the instruction-following ability of LLMs and guiding them with explicit prompts to explain their predictions (Figure 4). We call these prompting-based rationales.
@@
Figure 1: An example of our analysis methodology on the e-SNLI dataset. Human alignment compares model rationales with human-annotated rationale; Model faithfulness measures the rates when model prediction changes (e.g. from Contradiction to Entailment) after masking the rationales identified by the model. [^4] In the first prompt, we use the true label for human alignment, and the predicted label for faithfulness experiments.
@@
To evaluate different rationales, previous works on model interpretation establish two properties of rationales that are critical for successful interpretability: human alignment (DeYoung et al., 2019; Hase and Bansal, 2022) and faithfulness (Jacovi and Goldberg, 2020). Human alignment refers to the degree to which the rationales match or align with human-annotated rationales, while faithfulness assesses whether the rationales truly reflect the model’s internal process. A longstanding debate exists regarding the relationship between these two aspects (Agarwal et al., 2024). However, studies on LLM rationales either focus on the faithfulness of off-the-shelf LLMs (Huang et al., 2023a; Madsen et al., 2024), or their human alignment Chen et al. (2023), but lack a comprehensive exploration of the two properties together. Specifically, recent works (Huang et al., 2023a; Madsen et al., 2024) study prompting-based methods and show that they might not be faithful to the reasoning process of LLMs. Moreover, they only consider LLMs as out-of-box models, without fine-tuning on specific tasks. How fine-tuning of LLMs on downstream tasks influences the human alignment and faithfulness of LLM rationales is under-explored.
@@
In this paper, we conduct extensive experiments to evaluate LLM rationales more comprehensively and bridge the gap in existing research. We consider five state-of-the-art LLMs, encompassing both open-source models (Llama2 (Touvron et al., 2023), Llama3, Mistral (Jiang et al., 2023)) and proprietary models (GPT-3.5-Turbo, GPT-4-Turbo (OpenAI et al., 2024)). Our study leverages two annotated natural language classification datasets, e-SNLI (Camburu et al., 2018a) and MedicalBios (Eberle et al., 2023), to evaluate and compare rationale extraction methods based on prompting strategies and feature attribution-based techniques such as Input×Gradient (Li et al., 2016).
Through our experiments, we find that while prompting-based rationales are generally less faithful than attribution-based methods, they tend to align better with human-annotated rationales, both before and after fine-tuning. Surprisingly, even when prompting-based rationales exhibit poor performance, they can still produce explanations that reasonably align with human reasoning.
We also observe that low classification performance and collapsing predictions might be related to the faithfulness limitation of LLM rationales. Fine-tuning LLMs on specific datasets improves the quality of rationales for both prompting and, particularly, attribution techniques in terms of faithfulness and human alignment. This finding complements the observations in Madsen et al. (2024), where faithful evaluation was conducted using only out-of-the-box LLMs.
In summary, our work contributes to the ongoing efforts to enhance the interpretability and trustworthiness of LLMs by providing empirical evidence and practical recommendations for extracting and evaluating rationales from these models.
@@ 2 Related Work
Premise: Asian individuals are taking pictures of some unidentified thing.
Hypothesis: Asians play soccer in the park.
Label: Contradiction
[taking,
pictures,
play, soccer]
[Asian,
individuals, taking,
pictures, unidentified, thing, soccer, park]
[Asian,
individuals,
unidentified, soccer]
[Asian,
pictures,
play, soccer]
@@
training, MRCS]
[His, MRCS,
@@
[surgical, MRCS, SpR]
She works at both the Park Nicollet Women’s Center and Regions Hospital – Cancer Care Center. Kandice dedicates herself to providing nurturing, sensitive care along with good educational support to her patients.
Label: Nurse
[nurturing, sensitive, care] [Kandice, works,
nurse, Cancer Care Center, Women’s Center, provides, sensitive care, educational support, patients]
[care, patients, support] [care, patients]
Table 1: Examples from e-SNLI and MedicalBios datasets are presented, showcasing rationales from human annotations alongside those generated by three models (Using "Short Prompt Top-Var" method). Human rationales and the corresponding generated words are highlighted in bold font.
@@
Recent literature in natural language processing (NLP) has seen a surge in interpretability methods aimed at making models more transparent and understandable. The traditional interpretability methods include 1) attribution-based methods, which leverage the attention weights in models like transformers to identify which parts of the input the model focuses on when making a decision Vaswani et al. (2023); Clark et al. (2019); Abnar and Zuidema (2020), 2) Gradient-based methods, which provide explanations by identifying which input tokens most influence the model’s output, often using techniques like gradient-based saliency maps Simonyan et al. (2014a), or its extension by incorporating the input vector norms or integration Sundararajan et al. (2017). 3) Vector-based methods that propagate the decomposed representations throughout the model achieving the best faithfulness results on encoder-based models Kobayashi et al. (2020, 2021); Ferrando et al. (2022); Modarressi et al. (2022, 2023). More recently, researchers have been using rationales, written in natural language, to serve interpretability where they can reveal the “reasoning” behind model decisions.
@@ 3.1 Datasets
We utilize two natural language classification datasets that have been annotated with human rationales indicating which input words were pivotal for the ground truth label. Table 1 shows examples of these datasets alongside the human rationale annotation and model-generated rationale.
e-SNLI This dataset Camburu et al. (2018b) is a natural language inference task with 3 classes including Entailment, Contradiction, and Neutral, showing the relation between the premise and hypothesis sentences. This dataset is annotated for rationales supporting the classification label by DeYoung et al. (2019). We utilize 5,000 examples from the training set and 300 examples from the test set.
MedicalBios MedicalBios Eberle et al. (2023) consists of human rationale annotations for a subset of 100 samples (five medical classes) from the BIOS dataset De-Arteaga et al. (2019) for the occupation classification task.
Model Mistral-7B Instruct-v0.2 Llama-2-7b chat Llama-3-8B Instruct GPT-3.5 Turbo 1106 GPT-4 Turbo 2024-04-09
Dataset E-SNLI MedBios E-SNLI MedBios E-SNLI MedBios E-SNLI MedBios E-SNLI MedBios
Method Selection
Attribution-based
Attention Top-Ratio 28.59 29.34 40.92 22.11 23.80 15.01 - - - -
Saliency Top-Ratio 37.16 28.97 28.64 31.35 32.19 27.80 - - - -
Input×Gradient Top-Ratio 31.41 27.51 26.85 30.96 36.32 29.83 - - - -
Attention Top-Var 36.03 41.01 48.69 35.08 30.56 24.30 - - - -
Saliency Top-Var 46.07 38.30 37.25 41.49 39.78 36.36 - - - -
Input×Gradient Top-Var 38.86 37.08 36.66 40.45 44.76 39.48 - - - -
Prompting-based
Normal Prompt Unbound 38.95 42.50 45.45 49.16 39.58 49.05 43.96 45.14 47.24 51.89
Short Prompt Unbound 41.84 42.26 43.54 43.74 43.60 51.28 43.28 45.28 46.58 53.50
Normal Prompt Top-Ratio 41.17 41.17 45.90 43.98 35.98 47.58 36.58 43.68 45.27 47.81
Short Prompt Top-Ratio 41.83 42.05 40.91 45.65 37.46 46.91 37.92 41.00 47.31 45.64
Normal Prompt Top-Var 45.93 48.43 50.93 52.23 41.67 58.85 43.85 57.16 55.33 59.28
Short Prompt Top-Var 46.97 48.08 47.34 50.54 44.65 58.17 44.31 53.60 55.78 59.10
Table 2: Human alignment F1$\uparrow$ score in e-SNLI and MedicalBios datasets. Random baseline (Selecting Top-Var random words) is 27±4 and 22±1 for e-SNLI and MedicalBios respectively over 100 seeds. The top two alignments in each column are indicated by bold and underlined formatting.
@@ 3.2 Models
We employ five of the latest large language models, encompassing both open-source and proprietary ones. From the open-source models, we utilize Llama2 Touvron et al. (2023), LLama3, and Mistral Jiang et al. (2023). For proprietary models, we include GPT3.5-Turbo and GPT4-Turbo OpenAI et al. (2024). All models are prompted without sampling during generation, leading to deterministic outputs. You can see the exact model information in Table 6 in the appendix.
@@ 3.3.1 Prompting-Based Method
We employ various prompting strategies to explore the effects of prompt wording and model alignment in generating text of similar length to human-annotated rationale (Tables 12&13). The following prompts are used to evaluate these aspects.
We test two versions of prompts to examine how the clarity and length of the prompt influence the model’s output:
Normal Prompt This version provides a detailed explanation, including all the points the model should consider. It is longer and aims to ensure the model fully understands the task.
Short Prompt Given that long prompts may confuse LLMs, we also use a shorter version that conveys the necessary information in a few sentences.
We also experiment with three versions of the introduced prompts to manage the number of words the model generates.
Unbound Prompt In this method, the model is not restricted in the number of words it can generate. It needs to establish the appropriate length autonomously.
Top-Var Prompt This prompt requires the model to generate exactly the same number of words as in the human rationale annotations for each sentence. This method controls for word count in our experiments, enabling us to assess model alignment independently of its word importance threshold.
Top-Ratio Prompt
In this approach, the model is guided to identify the top $k$ most important words within a given sentence. The value of $k$ is derived from a predetermined percentage of the total word count in the input sentence, a ratio established based on the training set. For example, this ratio is set at 20% for sentences in the e-SNLI dataset and 13% for those in the MedicalBios dataset.
@@ 3.3.2 Attribution-Based Methods
We employ the Inseq library Sarti et al. (2023) to implement attribution-based methods for LLMs. Specifically, we select three available options: (i) Attention Weight Attribution, which utilizes the model’s internal attention weights Wiegreffe and Pinter (2019); (ii) Simple Gradients (Saliency), which is based on the gradients of the output with respect to the inputs Simonyan et al. (2014b); and (iii) Input×Gradient, which factors in both the input vector size and the gradient in its calculations Li et al. (2016). We choose these methods because of their demonstrated faithfulness in previous work on NLP models Atanasova et al. (2020); Modarressi et al. (2022, 2023), and their potential for efficient execution on large language models with limited computational resources.
@@ 4 Results
In this section, we delve into utilizing both prompting-based and attribution-based approaches to extract rationale from the model, focusing on two aspects: human alignment and model faithfulness. Furthermore, we conduct fine-tuning experiments on open LLMs to examine how this process influences alignment and faithfulness.

### v2

@@ Abstract
We study how well large language models (LLMs) explain their generations through rationales – a set of tokens extracted from the input text that reflect the decision-making process of LLMs. Specifically, we systematically study rationales derived using two approaches: (1) popular prompting-based methods, where prompts are used to guide LLMs in generating rationales, and (2) technical attribution-based methods, which leverage attention or gradients to identify important tokens. Our analysis spans three classification datasets with annotated rationales, encompassing tasks with varying performance levels. While prompting-based self-explanations are widely used, our study reveals that these explanations are not always as “aligned” with the human rationale as attribution-based explanations. Even more so, fine-tuning LLMs to enhance classification task accuracy does not enhance the alignment of prompting-based rationales. Still, it does considerably improve the alignment of attribution-based methods (e.g., Input×Gradient). More importantly, we show that prompting-based self-explanation is also less “faithful” than attribution-based explanations, failing to provide a reliable account of the model’s decision-making process. To evaluate faithfulness, unlike prior studies that excluded misclassified examples, we evaluate all instances and also examine the impact of fine-tuning and accuracy on alignment and faithfulness. Our findings suggest that inconclusive faithfulness results reported in earlier studies may stem from low classification accuracy. These findings underscore the importance of more rigorous and comprehensive evaluations of LLM rationales. [^1] Code and data will be released upon paper acceptance.
@@ 1 Introduction
The rise of large language models (LLMs) has significantly transformed the field of natural language processing (NLP) (Touvron et al., 2023; Team et al., 2023; OpenAI et al., 2024), enabling a wide range of applications from web question answering to complex reasoning tasks. However, if they cannot clearly and reliably explain their outputs (Ji et al., 2023), it limits their deployment in high-stakes scenarios.
@@
Figure 1: An example of our analysis methodology on the e-SNLI dataset. Human alignment compares model rationales with human-annotated rationale; Model faithfulness measures when model prediction changes (e.g. from Contradiction to Entailment) after masking the rationales identified by the model. [^3] In the first prompt, we use the true label for human alignment, and the predicted label for faithfulness experiments.
Rationales [^4] Also called self-explanation or extractive rationales in previous work (Huang et al., 2023a; Madsen et al., 2024)., i.e., tokens of the input text that are most influential to the models’ predictions, are widely studied in the NLP community prior to the era of LLMs to interpret model predictions (Lei et al., 2016; DeYoung et al., 2019; Wiegreffe and Pinter, 2019; Jacovi and Goldberg, 2020). For LLMs, besides attribution-based methods like attention weights (Wiegreffe and Pinter, 2019) or gradients (Li et al., 2016), rationales can also be extracted by leveraging the instruction-following ability of LLMs and guiding them with explicit prompts to explain their predictions (Figure 3). We call these prompting-based rationales.
@@
To evaluate different rationales, previous works on model interpretation establish two properties of rationales that are critical for successful interpretability: human alignment (DeYoung et al., 2019; Hase and Bansal, 2022) and faithfulness (Jacovi and Goldberg, 2020). Human alignment refers to the degree to which the rationales match or align with human-annotated rationales, while faithfulness assesses whether the rationales truly reflect the model’s internal process. However, studies on LLM rationales either focus on the faithfulness of off-the-shelf LLMs (Huang et al., 2023a; Madsen et al., 2024), or their human alignment Chen et al. (2023), but lack a comprehensive exploration of the two properties together. Moreover, they consider LLMs only as out-of-the-box models, without fine-tuning for specific tasks. The impact of fine-tuning to improve task accuracy on the alignment and faithfulness of LLM rationales remains under-explored.
@@
In this paper, we conduct extensive experiments to comprehensively evaluate LLM rationales and bridge the gap in existing research. We consider five state-of-the-art LLMs, encompassing both open-source models (Llama2 (Touvron et al., 2023), Llama3, Mistral (Jiang et al., 2023)) and proprietary models (GPT-3.5-Turbo, GPT-4-Turbo (OpenAI et al., 2024)). Our study leverages three annotated natural language classification datasets, e-SNLI (Camburu et al., 2018a), FEVER (Thorne et al., 2018a) and MedicalBios (Eberle et al., 2023), to evaluate and compare rationale extraction methods based on prompting strategies and feature attribution-based techniques such as Input×Gradient (Li et al., 2016).
Through our experiments, we find that attribution-based methods align more closely with humans in most cases, especially after fine-tuning. These methods also show more consistent improvements from fine-tuning compared to prompting. Additionally, we observe that low classification performance and collapsing predictions are linked to the limitations in evaluating the faithfulness of LLM rationales, shedding light on the task- and model-dependent findings of previous research. Most importantly we demonstrate that attribution-based methods offer a more faithful reflection of the model’s inner workings than prompting.
In summary, our work provides a systematic framework, empirical evidence, and practical recommendations for extracting and evaluating LLM rationales. Our findings highlight critical limitations in prompting-based rationales, emphasizing the need for further efforts to improve the interpretability and trustworthiness of LLMs.
@@ 2 Related Work
FEVER Claim: The Lincoln-Douglas debates happened in Alton, Illinois.
Evidence: The Lincoln – Douglas Debates -LRB- also known as The Great Debates of 1858 -RRB- were a series of seven debates between Abraham Lincoln … The debates were held in seven towns in the state of Illinois : Ottawa on August 21 … Alton, on, October, 15
Label: Supports
[Alton, on,
October, 15]
[Lincoln,
Douglas,
Illinois,
debates]
[Alton,
October,
Illinois,
Debates]
[Alton,
debates,
October,
15]
@@
training,
MRCS]
[His,
MRCS,
@@
[surgical, MRCS,
SpR]
Table 1: Examples from the datasets are presented, showcasing rationales from human annotations alongside those generated by three models. Human rationales and the corresponding generated words are highlighted in bold font.
@@
Recent literature in natural language processing (NLP) has seen a surge in interpretability methods aimed at making models more transparent and understandable. The traditional interpretability methods include 1) attention-based methods, which leverage the attention weights in models like transformers to identify which parts of the input the model focuses on when making a decision Vaswani et al. (2023); Clark et al. (2019); Abnar and Zuidema (2020), 2) Gradient-based methods, which provide explanations by identifying which input tokens most influence the model’s output, often using techniques like gradient-based saliency maps Simonyan et al. (2014a), or its extension by incorporating the input vector norms or integration Sundararajan et al. (2017). 3) Vector-based methods that propagate the decomposed representations throughout the model achieving the best faithfulness results on encoder-based models Kobayashi et al. (2020, 2021); Ferrando et al. (2022); Modarressi et al. (2022, 2023).
@@ 3.1 Datasets
We use three natural language classification datasets, each annotated with human rationales that highlight the key input words essential for determining the correct label. Table 1 shows examples of these datasets alongside the human rationale annotation and model-generated rationale.
e-SNLI This dataset Camburu et al. (2018b); DeYoung et al. (2019) is a natural language inference task with three classes including Entailment, Contradiction, and Neutral, showing the relation between the premise and hypothesis sentences. We utilize 5,000 examples from the training set and 300 examples from the test set.
FEVER The Fact Extraction and Verification dataset Thorne et al. (2018b); DeYoung et al. (2019) focuses on verifying claims based on provided evidence. Each claim is labeled as either "supports" or "refutes," with an annotated rationale indicating the specific portion of the evidence that underpins this classification. We use 5,000 samples from the training set and 300 from the test set.
MedicalBios Eberle et al. (2023) consists of human rationale annotations for a subset of 100 samples (five medical classes) from the BIOS dataset De-Arteaga et al. (2019) for the occupation classification task.
Figure 2: Classification accuracy throughout 10 epochs of fine-tuning. PT denotes the pre-trained model’s accuracy. "Zero Rate Baseline" refers to the performance of a classifier that assigns all inputs to the majority class. Our tasks include E-SNLI and FEVER, where pre-trained models tend to underperform, as well as MedicalBios, where they demonstrate strong performance.
@@ 3.2 Models
We employ five of the latest LLMs. From the open-source models, we utilize Llama2 Touvron et al. (2023), LLama3, and Mistral Jiang et al. (2023). For proprietary models, we include GPT3.5-Turbo and GPT4-Turbo OpenAI et al. (2024). All models are prompted without sampling during generation, leading to deterministic outputs (See Table A.1 for model information).
@@ 3.3.1 Prompting-Based Method
We employ different prompts for each stage of our experiments which are shown in Tables A.3, A.4, and A.5. Our prompts are engineered to convey the necessary information in a few sentences without confusing the models. We also experiment with two variations of each explanation prompt to manage the number of words the model generates.
Unbound Prompt In this method, the model autonomously determines the appropriate length of its generated text without word count restrictions.
Top-Var Prompt In this prompt, the model generates the top-$k$ words, where $k$ matches the exact number of words annotated in the human rationales for each sentence. This method controls for word count in our experiments, allowing us to evaluate model alignment independently of its word importance threshold and ensuring a fair evaluation of faithfulness by keeping the number of masked words consistent across experiments.
@@ 3.3.2 Attribution-Based Methods
We employ the Inseq library Sarti et al. (2023) to implement attribution-based methods for LLMs. Specifically, we select three available options: (i) Attention Weight Attribution, which utilizes the model’s internal attention weights Wiegreffe and Pinter (2019); (ii) Simple Gradients (Saliency), which is based on the gradients of the output with respect to the inputs Simonyan et al. (2014b); and (iii) Input×Gradient, which factors in both the input vector size and the gradient in its calculations Li et al. (2016). We choose these methods because of their demonstrated faithfulness in previous work on NLP models Atanasova et al. (2020); Modarressi et al. (2022, 2023), and their potential for efficient execution on large language models with limited computational resources. For each method, we focus on the output label word produced by the model in response to the classification prompt (Table A.3), calculating its attributions to the input tokens.
Pre-trained Model F1 Fine-tuned Model F1
E-SNLI FEVER MedBios E-SNLI FEVER MedicalBios
Model Method Selection
Mistral-7B Instruct-v0.2 Prompting Unbound 36.36 24.68 38.14 33.03 (−3.33) 25.64 (+0.95) 38.87 (+0.72)
Top-Var 40.08 26.24 44.11 35.29 (−4.79) 26.45 (+0.21) 45.02 (+0.92)
Attention Top-Var 36.26 37.23 37.36 42.14 (+5.87) 38.64 (+1.41) 38.29 (+0.93)
Saliency Top-Var 46.46 37.96 40.76 49.69 (+3.24) 40.44 (+2.47) 46.91 (+6.15)
Input×Grad Top-Var 40.45 36.36 40.10 42.70 (+2.25) 39.21 (+2.85) 45.38 (+5.28)
Llama-2-7b chat Prompting Unbound 38.57 14.69 32.64 38.44 (−0.13) 14.19 (−0.50) 33.87 (+1.22)
Top-Var 45.65 20.89 49.77 44.33 (−1.31) 20.43 (−0.46) 50.48 (+0.71)
Attention Top-Var 31.65 32.16 31.79 32.90 (+1.25) 32.94 (+0.78) 32.33 (+0.54)
Saliency Top-Var 34.92 30.52 36.49 46.56 (+11.64) 38.52 (+8.00) 34.12 (−2.37)
Input×Grad Top-Var 35.43 31.01 37.84 46.38 (+10.96) 38.05 (+7.03) 35.35 (−2.49)
Llama-3-8B Instruct Prompting Unbound 37.68 23.79 46.52 28.14 (−9.53) 21.83 (−1.96) 47.94 (+1.42)
Top-Var 43.07 28.76 59.99 30.03 (−13.04) 32.58 (+3.82) 59.76 (−0.23)
Attention Top-Var 28.67 33.87 26.48 28.60 (−0.07) 34.80 (+0.93) 29.57 (+3.10)
Saliency Top-Var 41.42 34.29 39.68 44.62 (+3.20) 35.61 (+1.32) 41.67 (+1.98)
Input×Grad Top-Var 44.59 35.27 45.61 49.22 (+4.63) 36.30 (+1.04) 46.68 (+1.07)
GPT-3.5 Turbo 1106 Prompting Unbound 43.14 22.76 42.95 - - -
Top-Var 46.06 34.27 53.96 - - -
GPT-4 Turbo 2024-04-09 Prompting Unbound 51.44 29.48 53.25 - - -
Top-Var 52.53 43.23 60.77 - - -
Random Random Top-Var 26.54 33.90 21.97 26.54 33.90 21.97
Table 2: Human alignment F1$\uparrow$ score. The difference in alignment between the fine-tuned (10 epochs) and pre-trained model is reported in the parentheses. Average random baseline (Selecting Top-Var random words) over 100 seeds is also reported. The highest alignment in each combination of model and dataset is underlined. Attribution-based methods are more aligned with humans in the majority of the cases, especially after fine-tuning. They also exhibit more consistent improvements by fine-tuning compared to prompting.
@@ 4 Results
In this section, we delve into utilizing both prompting-based and attribution-based approaches to extract rationale from the model, focusing on two aspects: human alignment and model faithfulness. Furthermore, we conduct fine-tuning experiments on open LLMs to examine how task performance influences alignment and faithfulness.
4.1 Task Performance
We begin our results by examining the accuracy of the models on the datasets. Figure 2 presents the pre-trained (PT) off-the-shelf accuracy of the models, along with their performance improvements during fine-tuning for 10 epochs.
As described by Wang et al. (2024) and Zhong et al. (2023), LLMs may underperform small fine-tuned models such as BERT, and smaller LLMs like LLaMA-2-7B might even collapse entirely. Madsen et al. (2024) reports similar behavior, with certain combinations of tasks and models performing poorly across different prompt variations. This pattern is evident in our results as well, where pre-trained open LLMs exhibit near-random accuracy on the e-SNLI and FEVER datasets (Figure 2).
Unlike previous work Madsen et al. (2024) that simply excluded misclassified examples from their faithfulness evaluations—which may represent more than half of the dataset—we include all examples in our analysis. Additionally, we fine-tune [^5] The hyperparameters are provided in Table A.2. the LLMs using LoRA Hu et al. (2022). As shown in Figure 2, the classification performance of LLaMA-2, LLaMA-3, and Mistral improves significantly after fine-tuning on the underperforming datasets, e-SNLI and FEVER, while also showing slight enhancements on the well-performing MedicalBios dataset. This fine-tuning narrows the performance gap with GPT-4-Turbo across all cases.
In the following sections, we will explore the alignment and faithfulness of pre-trained and fine-tuned models and investigate how task performance influences these aspects.